OpenStudy (anonymous):

2x-y>or=to 0. How do I do the y interslope form?

OpenStudy (gorv):

intercept slope form is y=mx+c 2x-y>=0 so move y on other side 2x>=y or y=<2x on comparing with y=mx+c we see that intercept c =0

OpenStudy (anonymous):

muchos gracias